OpenStudy (anonymous):

Find the equation of the line specified. The line passes through the points ( 2, 1) and ( 5, 4) a. y = x + 1 c. y = x - 1 b. y = -x - 1 d. y = x + 3

OpenStudy (texaschic101):

first you need to find the slope using the slope formula... slope(m) = (y2 - y1) / (x2 - x1) (2,1) x1 = 2 and y1 = 1 (5,4) x2 = 5 and y2 = 4 now we sub slope(m) = (4 - 1) / (5 - 2) can you finish and tell me the slope ?

OpenStudy (anonymous):

whats the answer im on a time crunch

OpenStudy (texaschic101):

slope is 1. y = mx + b slope(m) = 1 (2,1) x = 2 and y = 1 now sub 1 = 1(2) + b 1 = 2 + b 1 - 2 = b -1 = b equation is : y = -x - 1